1 friend leaves his house walking 3 mph. The other friend leaves his house walking 2 mph. Both are walking toward each other. Their houses are 4 miles apart. How long will it be before they meet each other?
Here is what I have tried:
4 miles= 3r+ 2r/2

Answer by mananth(16946)

1 friend leaves his house walking 3 mph. The other friend leaves his house walking 2 mph. Both are walking toward each other. Their houses are 4 miles apart. How long will it be before they meet each other?
1. speed = 3 mph
2. speed = 2 mph
...
walking towards each other.
speed = 2+3 = 5 mph.
distance = 4 miles
..
toime = d/r
t=4/3
t= 4/3 *60
t= 80 minutes =
1 hour 20 minutes
...
m.ananth@hotmail.ca